OpenStudy (ray1998):

What are the solutions of the quadratic equation? 2x^2 - 16x + 32 = 0 Will give fan and medal. Thank you in advance! :)

OpenStudy (anonymous):

The solution is x = 4

OpenStudy (ray1998):

Okay. Um, thank you, but could you tell me how? Like walk through with me how you got that answer? @Cardinal_Carlo

OpenStudy (anonymous):

Do you know the quadratic formula?

OpenStudy (ray1998):

No, sorry. @OrangeMaster

OpenStudy (anonymous):

\[2(x ^{2} - 8x + 16) = 0\] \[(x-4)^{2} = 0\] \[x = 4\]

OpenStudy (anonymous):

It is essential to answering the question. It is x= -b +/- sqrt(b^2-4ac)/2a

OpenStudy (anonymous):

The quadratic formula is too complicated and therefore impractical to use it in this particular case. But it is similarly effective.

OpenStudy (anonymous):

\[x=\frac{ -b \pm \sqrt{b^{2}-4ac} }{ 2a }\]

OpenStudy (ray1998):

Okay. How do you apply that to the equation though?

OpenStudy (ray1998):

It's not needed?

OpenStudy (anonymous):

2x^2 - 16x + 32 = 0 The 2 in front of x is 'a' -16 is 'b' 32 is 'c' Input the values and solve.

OpenStudy (anonymous):

\[x=\frac{ -(-16) \pm \sqrt{-16^{2}-4(2)(32)}}{ 2(2) }\]

OpenStudy (ybarrap):

Here's a fun way $$ (x - ?)^2+16 - ?^2=0\\ $$ To find that ? just take the coefficient of x, in this case 8 and divide it by 2 $$ (x - \cfrac{8}{2})^2+16 - \left (\cfrac{8}{2}\right )^2=0\\ $$ Just another approach.

OpenStudy (anonymous):

-(-16)=16 -16^2=256 4*2*32=256 256-256=0 sqrt(0)=0 2*2=4 16/4=4

OpenStudy (anonymous):

The plus/minus has no point when the sqrt of 4ac is 0

OpenStudy (ray1998):

-3^2 + 16 - 4^2 = -9 ? @ybarrap

OpenStudy (anonymous):

@ybarrap method works but it's confusing for starters

OpenStudy (anonymous):

So, @Cardinal_Carlo is correct, but there's the explanation.

OpenStudy (anonymous):

Or, an explanation rather.

OpenStudy (ray1998):

Okay. Gotcha. Sorry guys, let me just read through what all of you said :) thank you very very much for helping me. I'm trying to understand. @OrangeMaster @Cardinal_Carlo @ybarrap

OpenStudy (anonymous):

No problem, just try to use what is easiest for you to understand. :)

OpenStudy (ybarrap):

No x variables? $$ -3^2 + 16 - 4^2 = -9 $$ If so $$ -3^2 + 16 - 4^2 =-9+16-16=-9 $$ So the statement is true.

OpenStudy (ray1998):

I think I'm starting to understand it a little. I see both of your methods and I feel a better about them :)

OpenStudy (anonymous):

The quadratic formula may be hard to memorize and understand at first, but it will get the problem solved.

OpenStudy (ray1998):

Okay. Thank you.

OpenStudy (ray1998):

Thank you guys SO much! @OrangeMaster @Cardinal_Carlo @ybarrap

OpenStudy (ybarrap):

You're welcome $$ \color{green}{\Huge {:-)}} $$

OpenStudy (anonymous):

You're welcome